# Bounded geometry for PCF-special subvarieties

### v1

@@ 1. Introduction
In this article, we study the geometry of the set $\mathrm{PCF}$ by examining how it intersects with families of varieties in $\mathrm{M}_{d}$, and we prove uniform bounds on the degrees of PCF-special subvarieties in these families. Here, and throughout this article, we measure degree with respect to a fixed choice of compactification $\mathrm{M}_{d}$ (as defined in [Si1]) and a very ample line bundle on $\mathrm{M}_{d}$; the degree bounds will depend on these choices, but the following result holds for any such choice:
@@
Fix integers $d\geq 2$ and and $D\geq 1$. There exist constants $B=B(D,d)$ and $N=N(D,d)$ so that for any complex algebraic subvariety $X\subset\mathrm{M}_{d}$ of degree $\leq D$, the Zariski closure
@@
Consider the curves $\mathrm{Per}_{1}(\lambda)$ in $\mathrm{M}_{2}$, parameterized by $\lambda\in\mathbb{C}$, consisting of all quadratic maps having a fixed point of multiplier $\lambda$. This family was first introduced by Milnor in [Mi1], and they are lines in the natural coordinates that Milnor introduced on $\mathrm{M}_{2}\simeq\mathbb{C}^{2}$. It is known that $\mathrm{Per}_{1}(\lambda)$ is PCF-special if and only if $\lambda=0$ [DWY]. And although $\mathrm{PCF}$ is Zariski-dense in $\mathrm{M}_{2}$, Theorem 1.1 combined with the main result of [DWY] implies: There is a uniform bound $M$ so that
@@
In fact we prove stronger versions of Theorem 1.1 and Theorem 1.2 in the spirit of the Bogomolov Conjecture [Ul, Zh1] or, more recently, of the “height gap principle” [GGK, Gao] in the setting of abelian varieties, for subvarieties of $\mathrm{M}_{d}$ defined over $\overline{\mathbb{Q}}$. We work with the intersection theory for adelically-metrized line bundles on quasiprojective varieties, as developed by Yuan and Zhang [YZ]. The critical height of a conjugacy class $[f]\in\mathrm{M}_{d}(\overline{\mathbb{Q}})$ is defined by
@@
To state our results, we fix a projective embedding of $\mathrm{M}_{d}$ defined over $\mathbb{Q}$, allowing us to define degrees of subvarieties. We also fix height functions $h_{\mathrm{Ch}}$ on the associated Chow varieties $\mathrm{Ch}(\mathrm{M}_{d},r,D)$ of cycles of dimension $r$ and degree $\leq D$ in $\mathrm{M}_{d}$, for example a Philippon height [Ph]; see §2.5 for definitions. The constants $B$, $N$, and $\epsilon$ of the following theorems will depend on these choices.
@@
Fix integers $d\geq 2$ and and $D\geq 1$. There exist constants $B=B(D,d)$, $N=N(D,d)$, and $\epsilon=\epsilon(D,d)>0$ so that for any irreducible algebraic subvariety $X\subset\mathrm{M}_{d}$ of degree $\leq D$ defined over $\overline{\mathbb{Q}}$, and every $0\leq\epsilon^{\prime}<\epsilon$, the Zariski closure
@@ 1.1. History and context
The study of the PCF-special varieties in $\mathrm{M}_{d}$ grew out of parallels between the theory of complex dynamical systems and the study of abelian varieties. A conjectural classification of PCF-special subvarieties in $\mathrm{M}_{d}$ first appears in [Si2, Chapter 6] where it was given the name of “Dynamical André-Oort” (or “DAO” as it is known in [JX]), because the geometry of the PCF locus in $\mathrm{M}_{d}$ shares features with the locus of CM points in a moduli space of abelian varieties; see [Za] for background on more general Unlikely Intersection problems. The first cases of DAO were proved by Baker-DeMarco [BD] and Ghioca-Hsia-Tucker [GHT], followed by a series of results mostly focused on polynomial dynamics where the relations and symmetries are understood, such as [GY, FG2], culminating in a complete classification of PCF-special curves in moduli spaces of polynomials by Favre-Gauthier [FG3] and a characterization of PCF-special curves in $\mathrm{M}_{d}$ by Ji-Xie [JX].
@@
We point out that the aforementioned uniformity results each relied crucially on the classification of the ‘special’ varieties in their respective settings. We do not know of any result in the literature where the uniformity is known without a classication. In contrast, here, we do not prove that PCF-special subvarieties are necessarily defined by critical orbit relations, as conjectured in [BD]. In particular, we do not recover (or use) the main result of [JX] for curves in $\mathrm{M}_{d}$, but each of our theorems is new also for curves in $\mathrm{M}_{d}$.
@@ 1.2. Proof strategy
To prove Theorem 1.7, we first pass to a maximally varying family of varieties $\mathcal{X}^{\prime}\to V$ in a finite branched cover $\mathcal{M}_{d}^{cm}\to\mathrm{M}_{d}$ that parameterizes maps on $\mathbb{P}^{1}$ with marked critical points $c_{1},\ldots,c_{2d-2}:\mathcal{M}_{d}^{cm}\to\mathbb{P}^{1}$. We construct two nonzero bifurcation measures on the fiber power $(\mathcal{X}^{\prime})_{V}^{\ell+1}$, from two distinct $r$-tuples of critical points, and, associated to these measures, we define two adelically metrized line bundles on the quasiprojective $(\mathcal{X}^{\prime})_{V}^{\ell+1}$ over $\mathbb{Q}$, in the sense of [YZ]. The construction of the measures is carried out in Section 3, and their nonvanishing ultimately comes from properties of the bifurcation current and its powers in $\mathrm{M}_{d}$ from [Mc1, De1, Du, GOV]. The nonvanishing of the measures implies that the two metrized line bundles are non-degenerate (as defined in [YZ]). We are thus able to deduce from the height inequality of [YZ, Theorem 1.3.2] that, if $\epsilon>0$ is sufficiently small, then “most” points of $Z_{\epsilon}(\mathcal{X})$ lie in subvarieties $X_{\lambda}$ of bounded $h_{\mathrm{Ch}}$-height.
@@
The contradiction will come from an analysis of the two measures. More precisely, using a slicing argument as in [MaS] (which was revisited in [DM1]), it suffices to analyze the family of bifurcation measures they induce on the the fibers $X_{\lambda}$. Here we appeal to the maximal variation of the family and the dynamical theory of $J$-stability. We find a parameter $\lambda_{0}\in V$ and conjugacy class $[f_{0}]\in X_{\lambda_{0}}$ so that one collection of $r$ critical points is pre-repelling for $f_{0}$, while one of the critical points in the other $r$-tuple is in the basin of a parabolic point. We then zoom into the parameter space at $f_{0}$, where the first $r$-tuple of preperiodic critical points yields an asymptotic similarity between the bifurcation measure and the dynamical space, exactly as in the proof of [BE, Main Theorem]. But zooming in at this rate yields 0 for the bifurcation measure defined with the second $r$-tuple of critical points, because one of the critical points has infinite orbit and is attracted to a parabolic cycle. Details are provided in Section 4. The proofs of Theorems 1.1, 1.2, 1.5, and 1.6 are completed in Section 5.
@@ 1.3. Questions
The question has only been answered in a few special cases with $r=1$: (1) for the 1-parameter families of polynomials treated by Baker-DeMarco in [BD] parameterized by $\Lambda=\mathbb{C}$, and (2) for the curves $\mathrm{Per}_{1}(\eta)$ in $\mathrm{M}_{2}$, for each $\eta\in\mathbb{C}$, of Example 1.4 [DWY].
Question 1.9.
Suppose that $\Lambda$ is a smooth and irreducible, quasiprojective, complex algebraic variety of dimension $r\geq 2$ and that $f:\Lambda\times\mathbb{P}^{1}\to\Lambda\times\mathbb{P}^{1}$ is a morphism that defines an algebraic family of maps on $\mathbb{P}^{1}$, with marked critical points. With the notation of Question 1.8, suppose that
$T_{1}\wedge\cdots\wedge T_{r}\not=0\quad\mbox{ and }\quad T_{r+1}\not=0$
@@
on $\Lambda$. Can we conclude that there exists $i\in\{1,\ldots,r\}$ so that the pair $(c_{i},c_{r+1})$ is dynamically related along $\Lambda$? (We remark here that one can define a dynamically related tuple $(c_{1},\ldots,c_{r+1})$ similarly; see [De2, §6.2]. But it follows by [MeSc, Proposition 2.21], that dynamical relations always come in pairs.)
@@ 1.4. Acknowledgements
We would like to thank Dori Bejleri, Xavier Buff, Ziyang Gao, Thomas Gauthier, Zhuchao Ji, Lars Kühne, Harry Schmidt, Gabriel Vigny, and Junyi Xie for helpful discussions. Our research was supported in part by the National Science Foundation. In addition, Laura acknowledges the Radcliffe Institute for Advanced Study at Harvard; Myrto acknowledges the University of Toronto and Fields Insitute; and Hexi acknowledges Zhejiang University for their support of this collaboration.
@@ 2.4. The critical height
defines a function on $\mathcal{M}_{d}^{cm}$, which Silverman called the critical height and conjectured that it induces a Weil height on $\mathrm{M}_{d}\setminus\mathcal{L}_{d}$; see, for example, [Si2]. This was proved by Ingram in [In].
The construction of the critical height was revisited by Yuan and Zhang in [YZ, §6.3], where they show that $\hat{h}_{\mathrm{crit}}$ is the height associated to a non-degenerate adelically metrized line bundle on $\mathcal{M}^{cm}_{d}$ over $\mathbb{Q}$. Indeed, they construct an $f$-invariant adelic line bundle $\overline{L}_{f}$ on $\mathcal{M}_{d}^{cm}\times\mathbb{P}^{1}$. The critical height is then the height function associated to the ‘restriction’ of $\overline{L}_{f}$ to the ramification divisor of $f$; see [YZ, §6.3]. Important for us is that the curvature distribution at the archimedean place is equal to the bifurcation current $T_{\mathrm{bif}}$ defined in §2.3.
@@ 2.5. The Chow variety and maximal variation
Suppose that $Y$ is an irreducible quasiprojective algebraic variety, equipped with a projective embedding of $Y\hookrightarrow\mathbb{P}^{N}$ defined over $\mathbb{Q}$. We let $\overline{Y}$ denote the closure of $Y$ in $\mathbb{P}^{N}$. For each integer $0\leq r<\dim Y$, and with respect to this embedding, we first consider the Chow variety $\mathrm{Ch}(\overline{Y},r,D)$ of algebraic cycles $X\subset\overline{Y}$ in dimension $r$ with degree $\leq D$. It is an algebraic subvariety, usually reducible, of the Chow variety $\mathrm{Ch}(\mathbb{P}^{N},r,D)$, consisting of the cycles with dimension $r$ and degree $\leq D$ supported in $\overline{Y}$. As $\overline{Y}$ is defined over $\mathbb{Q}$, the Chow variety $\mathrm{Ch}(\overline{Y},r,D)$ is also defined over $\mathbb{Q}$. We let $\mathrm{Ch}(Y,r,D)$ denote the Zariski-open subset consisting of cycles with support intersecting $Y$.
@@ 3. Bifurcation measures on $\mathcal{X}_{V}^{\ell}$ and $\mathcal{X}_{V}^{\ell+1}$
When the marked points are critical points, with $f:=f_{1}=\cdots=f_{m}$, Dujardin proved that there is a dense set of parameters in $\operatorname{supp}T_{1}\wedge\cdots\wedge T_{r}$ in $\mathcal{M}_{d}^{cm}$ at which the critical point ${\bf c}=(c_{1},\ldots,c_{r})$ for $(f,\ldots,f)$ is transversely prerepelling over $\mathcal{M}_{d}^{cm}$ [Du, Theorem 1.1]. And conversely, if the $r$-tuple of critical points is properly prerepelling for $(f,\ldots,f)$ at some parameter over a parameter space of dimension $\geq r$, then $T_{1}\wedge\cdots\wedge T_{r}$ must be nonzero in that parameter space, as first observed in [BE] (where it was proved in the case $r=2d-2$ and assuming transversely prerepelling) and revisited in [Ga1, Theorem 6.2] and [BBD, Proposition 3.7] in more general contexts and without assuming transversality. See [DM1, Proposition 4.8] for the version we use, because it covers the setting we need.
@@ 3.1. Measure on $\mathcal{X}_{V}^{\ell}$
If ${\bf c}$ satisfies the conclusion of Proposition 3.1 for $\Phi$, will say that it witnesses the nonvanishing of $\mu_{\ell}$.
@@ 3.2. Measures on $\mathcal{X}_{V}^{\ell+1}$
In what follows it is convenient to work with the Deligne pairing (of relative dimension dimension 00), of the invariant metrized line bundles restricted to graphs of critical marked points (see [YZ, §4.2] for metrics of the Deligne pairing and [YZ, §6.2.2] for the ‘restriction’ of the invariant metrized line bundle to a subvariety). By [YZ, Theorem 4.2.3], to each marked tuple of critical points ${\bf c}:\mathcal{X}^{\ell}_{V}\to(\mathbb{P}^{1})^{\dim\mathcal{X}^{\ell}_{V}}$ we can associate a metrized line bundle $\langle\overline{L}_{\Phi}|_{\Gamma_{\bf c}}\rangle\in\widehat{\mathrm{Pic}}(\mathcal{X}^{\ell}_{V})_{\mathbb{Q},\mathrm{nef}}$, where $\Gamma_{\bf c}$ denotes its graph in $\mathcal{X}^{\ell}_{V}\times(\mathbb{P}^{1})^{\dim\mathcal{X}^{\ell}_{V}}$. Moreover, we let
@@
and note that their curvature forms are the bifurcation currents
@@
Recall that $T_{f_{i},c_{j}}^{\wedge 2}=0$ for all $i,j$ and similarly $(\Pi^{*}_{2}c_{1}(\overline{L}_{i}))^{\wedge r+1}=(\Pi_{2})^{*}(c_{1}(\overline{L}_{i})^{\wedge r+1})=0$ for $i=1,2$, by the definition of the bifurcation current as limits of pullbacks, because the map $\Phi_{(\ell+1)}$ acts on the $r$-dimensional $(\mathbb{P}^{1})^{r}$. Moreover, since $\dim\mathcal{X}^{\ell}_{V}=\ell(r+1)$ we also have that $\Pi_{1}^{*}(c_{1}(\overline{L}))^{\wedge\ell(r+1)+1}=0$. Thus
@@
for all $(\lambda,f_{1},\ldots,f_{\ell+1})\in U(\overline{\mathbb{Q}})$ for a Zariski open and dense subset $U\subset\mathcal{X}^{\ell+1}_{V}$. It remains to show that for some $\epsilon=\epsilon(\mathcal{X})$, the set
@@
as measures on $X_{\lambda}$ for $\pi_{*}\mu_{\ell}$-almost every parameter $\lambda\in V(\mathbb{C})$. Indeed, suppose there exists $b\in V(\mathbb{C})$ where (4.8) doesn’t hold, and so that $\pi_{*}\mu_{\ell}(U)>0$ for every open neighborhood $U$ of $b$. Then we can find a continuous, non-negative, compactly supported function $\psi$ on $\mathcal{X}$ such that
@@
We conclude from (4.8) that
(4.9) $\mu_{\lambda}^{r+1}=\alpha\,\mu_{\lambda}^{1}$
for all $\lambda$ in the support of $\pi_{*}\mu_{\ell}$ in $V$, by continuity of the measures.
Recall from Proposition 3.2 that we have parameter $x_{1}=(\lambda_{1},s_{1},\ldots,s_{\ell})\in\operatorname{supp}\mu_{\ell}$ so that the tuple of marked critical points $(c_{1},\ldots,c_{r})$ is transversely prerepelling for $(f_{\ell},\ldots,f_{\ell})$ at $s_{\ell}$ along $X_{\lambda_{1}}$. This implies that $\mu_{\lambda_{1}}^{r+1}$ is nonzero. Moreover, exactly as in [BE, Section 3] and the generalization appearing in [Ga2, Proposition 3.2], we know that if we zoom in on the measure $\mu^{r+1}_{\lambda_{1}}$ on $X_{\lambda_{1}}$ at the parameter $s_{\ell}$ at an exponential rate in $n$ (determined by the multipliers of the repelling cycles in the forward orbits of the marked critical points), and rescale the measure by $d^{nr}$, we get a nontrivial limit.
Let us now recall from Proposition 3.2 that the critical point $c_{r+1}$ is attracted to a parabolic cycle for $f_{\ell}$ at this parameter $x_{1}$. By (4.9), we also have that $\mu_{\lambda_{1}}^{1}=\alpha^{-1}\mu_{\lambda_{1}}^{r+1}\not=0$ on $X_{\lambda_{1}}$. But because of the parabolic point, we claim that if we zoom into the measure $\mu_{\lambda_{1}}^{1}$ on $X_{\lambda_{1}}$ at $s_{\ell}$ at the same exponential rate (which is supposed to give a nonzero limit), we actually get 0. This will be a contradiction.
@@
Suppose $f_{x}$ is a holomorphic family of maps on $\mathbb{P}^{1}$ parametrized by $x\in\mathbb{D}^{r}\subset\mathbb{C}^{r}$, with a marked critical point $c:\mathbb{D}^{r}\to\mathbb{P}^{1}$. Assume that $c(0)$ attracted to a parabolic fixed point at $z=0$ for $f_{0}$. Then for any constant $\rho>1$, there exists $\delta>0$ so that the sequence of functions

### v2

@@ 1. Introduction
In this article, we study the geometry of the set $\mathrm{PCF}$ by examining how it intersects with families of varieties in $\mathrm{M}_{d}$, and we prove uniform bounds on the degrees of PCF-special subvarieties in these families. Here, and throughout this article, we measure degree with respect to a fixed choice of very ample line bundle on $\mathrm{M}_{d}$, for example providing the projective compactification of $\mathrm{M}_{d}$ defined in [Si1]. The degree bounds will depend on the choice, but the following result holds for any such choice:
@@
Fix integers $d\geq 2$ and $D\geq 1$. There exist constants $B=B(D,d)$ and $N=N(D,d)$ so that for any complex algebraic subvariety $X\subset\mathrm{M}_{d}$ of degree $\leq D$, the Zariski closure
@@
Consider the family of curves $\mathrm{Per}_{1}(\lambda)$ in $\mathrm{M}_{2}$, parameterized by $\lambda\in\mathbb{C}$, consisting of all quadratic maps having a fixed point of multiplier $\lambda$. This family was first introduced by Milnor in [Mi1], and they are lines in the natural coordinates that Milnor introduced on $\mathrm{M}_{2}\simeq\mathbb{C}^{2}$. It is known that $\mathrm{Per}_{1}(\lambda)$ is PCF-special if and only if $\lambda=0$ [DWY]. And although $\mathrm{PCF}$ is Zariski-dense in $\mathrm{M}_{2}$, Theorem 1.1 combined with the main result of [DWY] implies: There is a uniform bound $M$ so that
@@
In fact we prove stronger versions of Theorem 1.1 and Theorem 1.2 in the spirit of the Bogomolov Conjecture [Ul, Zh1] or, more recently, of the “height gap principle” [GGK, Gao] in the setting of abelian varieties, for subvarieties of $\mathrm{M}_{d}$ defined over $\overline{\mathbb{Q}}$. The critical height of a conjugacy class $[f]\in\mathrm{M}_{d}(\overline{\mathbb{Q}})$ is defined by
@@
To state our results, we fix a projective embedding of $\mathrm{M}_{d}$ defined over $\mathbb{Q}$. We also fix height functions $h_{\mathrm{Ch}}$ on the associated Chow varieties $\mathrm{Ch}(\mathrm{M}_{d},r,D)$ of cycles of dimension $r$ and degree $\leq D$ in $\mathrm{M}_{d}$, for example a Philippon height [Ph]; see §2.5 for definitions. The constants $B$, $N$, and $\epsilon$ of the following theorems will depend on these choices.
@@
Fix integers $d\geq 2$ and $D\geq 1$. There exist constants $B=B(D,d)$, $N=N(D,d)$, and $\epsilon=\epsilon(D,d)>0$ so that for any irreducible algebraic subvariety $X\subset\mathrm{M}_{d}$ of degree $\leq D$ defined over $\overline{\mathbb{Q}}$, and every $0\leq\epsilon^{\prime}<\epsilon$, the Zariski closure
@@ 1.1. History and context
The study of the PCF-special varieties in $\mathrm{M}_{d}$ grew out of parallels between the theory of complex dynamical systems and the study of abelian varieties. A conjectural classification of PCF-special subvarieties in $\mathrm{M}_{d}$ first appears in [Si2, Chapter 6] where it was given the name of “Dynamical André-Oort” (or “DAO” as it is known in [JX]), because the geometry of the PCF locus in $\mathrm{M}_{d}$ shares features with the locus of CM points in a moduli space of abelian varieties; see [Za] for background on several Unlikely Intersection problems. The first cases of DAO were proved by Baker-DeMarco [BD] and Ghioca-Hsia-Tucker [GHT], followed by a series of results mostly focused on polynomial dynamics where the relations and symmetries are understood, such as [GY, FG2], culminating in a complete classification of PCF-special curves in moduli spaces of polynomials by Favre-Gauthier [FG3] and a characterization of PCF-special curves in $\mathrm{M}_{d}$ by Ji-Xie [JX]. We note, however, that there are currently no known classification results for PCF-special subvarieties of dimensions $>1$, even when restricting to spaces of polynomials. Nevertheless, Theorem 1.1 can be viewed as a “Uniform DAO” that holds for PCF-special subvarieties of any dimension.
@@
Remark 1.7.
We emphasize that the aforementioned uniformity results each relied crucially on the classification of the ‘special’ varieties in their respective settings. In contrast, the classification of PCF-special subvarieties is only conjectural [BD]; we obtain uniformity without proving that PCF-special subvarieties are necessarily defined by critical orbit relations. In particular, we do not recover (or use) the main result of [JX] for curves in $\mathrm{M}_{d}$, and each of our theorems is new also for curves in $\mathrm{M}_{d}$. We do not know of any other result in the literature where uniformity is known without the classication.
@@ 1.2. Proof strategy
To prove Theorem 1.8, we first pass to a maximally varying family of varieties $\mathcal{X}^{\prime}\to V$ in a finite branched cover $\mathcal{M}_{d}^{cm}\to\mathrm{M}_{d}$ that parameterizes maps on $\mathbb{P}^{1}$ with marked critical points $c_{1},\ldots,c_{2d-2}:\mathcal{M}_{d}^{cm}\to\mathbb{P}^{1}$. We construct two nonzero bifurcation measures on the fiber power $(\mathcal{X}^{\prime})_{V}^{\ell+1}$, from two distinct $r$-tuples of critical points, and, associated to these measures, we define two adelically metrized line bundles on the quasiprojective $(\mathcal{X}^{\prime})_{V}^{\ell+1}$, in the sense of [YZ]. The construction of the measures is carried out in Section 3, and their nonvanishing ultimately comes from properties of the bifurcation current and its powers in $\mathrm{M}_{d}$ from [Mc1, De1, Du, GOV]. The nonvanishing of the measures implies that the two metrized line bundles are non-degenerate (as defined in [YZ]). We are thus able to deduce from the height inequality of [YZ, Theorem 1.3.2] that, if $\epsilon>0$ is sufficiently small, then “most” points of $Z_{\epsilon}(\mathcal{X})$ lie in subvarieties $X_{\lambda}$ of bounded $h_{\mathrm{Ch}}$-height.
@@
The contradiction will come from an analysis of the two measures. More precisely, using a slicing argument as in [MaS] (which was revisited in [DM1]), it suffices to analyze the family of bifurcation measures they induce on the fibers $X_{\lambda}$. Here we appeal to the maximal variation of the family and the dynamical theory of $J$-stability. We find a parameter $\lambda_{0}\in V$ and conjugacy class $[f_{0}]\in X_{\lambda_{0}}$ so that one collection of $r$ critical points is pre-repelling for $f_{0}$, while one of the critical points in the other $r$-tuple is in the basin of a parabolic point. We then zoom into the parameter space at $f_{0}$, where the first $r$-tuple of preperiodic critical points yields an asymptotic similarity between the bifurcation measure and the dynamical space, exactly as in the proof of [BE, Main Theorem]. But zooming in at this rate yields 0 for the bifurcation measure defined with the second $r$-tuple of critical points, because one of the critical points has infinite orbit and is attracted to a parabolic cycle. Details are provided in Section 4. The proofs of Theorems 1.1, 1.2, 1.5, and 1.6 are completed in Section 5.
@@ 1.3. Questions
The question has only been answered in a few special cases with $r=1$: (1) for the 1-parameter families of polynomials treated by Baker-DeMarco in [BD] parameterized by $\Lambda=\mathbb{C}$, and (2) for the curves $\Lambda=\mathrm{Per}_{1}(\eta)$ in $\mathrm{M}_{2}$, for each $\eta\in\mathbb{C}$, of Example 1.4 [DWY].
Question 1.10.
Suppose that $\Lambda$ is a smooth and irreducible, quasiprojective, complex algebraic variety of dimension $r\geq 2$ and that $f:\Lambda\times\mathbb{P}^{1}\to\Lambda\times\mathbb{P}^{1}$ is a morphism that defines an algebraic family of maps on $\mathbb{P}^{1}$, with marked critical points. With the notation of Question 1.9, suppose that
$T_{1}\wedge\cdots\wedge T_{r}\not=0$
@@
on $\Lambda$. Can we conclude that there exists $i\in\{1,\ldots,r-1\}$ so that the pair $(c_{i},c_{r+1})$ is dynamically related along $\Lambda$? (We remark here that one can define a dynamically related tuple $(c_{1},\ldots,c_{r+1})$ similarly; see [De2, §6.2]. But it follows by [MeSc, Proposition 2.21], that dynamical relations always come in pairs.)
@@ 1.4. Acknowledgements
We would like to thank Dori Bejleri, Xavier Buff, Ziyang Gao, Thomas Gauthier, Zhuchao Ji, Lars Kühne, Harry Schmidt, Gabriel Vigny, and Junyi Xie for helpful discussions. Our research was supported in part by the National Science Foundation and by the Natural Sciences and Engineering Research Council of Canada. In addition, Laura acknowledges the Radcliffe Institute for Advanced Study at Harvard; Myrto acknowledges the University of Toronto and Fields Insitute; and Hexi acknowledges Zhejiang University for their support of this collaboration.
@@ 2.4. The critical height
defines a function on $\mathcal{M}_{d}^{cm}$, which Silverman called the critical height and conjectured that it is comparable with a Weil height on $\mathrm{M}_{d}\setminus\mathcal{L}_{d}$; see, for example, [Si2]. This was proved by Ingram in [In].
The construction of the critical height was revisited by Yuan and Zhang in [YZ, §6.3], where they show that $\hat{h}_{\mathrm{crit}}$ is the height associated to a non-degenerate adelically metrized line bundle on $\mathcal{M}^{cm}_{d}$ over $\mathbb{Q}$. Indeed, they construct an $f$-invariant adelic line bundle $\overline{L}_{f}$ on $\mathcal{M}_{d}^{cm}\times\mathbb{P}^{1}$ for the $f$ of (2.1). The critical height is then the height function associated to the ‘restriction’ of $\overline{L}_{f}$ to the ramification divisor of $f$; see [YZ, §6.3]. Important for us is that the curvature distribution at the archimedean place is equal to the bifurcation current $T_{\mathrm{bif}}$ defined in §2.3. The non-degeneracy of this metrized line bundle follows because the bifurcation measure is non-trivial, as was proved in [BB].
@@ 2.5. The Chow variety and maximal variation
Suppose that $Y$ is an irreducible quasiprojective algebraic variety, equipped with a projective embedding of $Y\hookrightarrow\mathbb{P}^{N}$ defined over $\mathbb{Q}$. We let $\overline{Y}$ denote the closure of $Y$ in $\mathbb{P}^{N}$. For each integer $0\leq r<\dim Y$, and with respect to this embedding, we first consider the Chow variety $\mathrm{Ch}(\overline{Y},r,D)$ of algebraic cycles $X\subset\overline{Y}$ in dimension $r$ with degree $\leq D$. It is an algebraic subvariety, possibly reducible, of the Chow variety $\mathrm{Ch}(\mathbb{P}^{N},r,D)$, consisting of the cycles with dimension $r$ and degree $\leq D$ supported in $\overline{Y}$. As $\overline{Y}$ is defined over $\mathbb{Q}$, the Chow variety $\mathrm{Ch}(\overline{Y},r,D)$ is also defined over $\mathbb{Q}$. We let $\mathrm{Ch}(Y,r,D)$ denote the Zariski-open subset consisting of cycles with support intersecting $Y$.
@@ 3. Bifurcation measures on $\mathcal{X}_{V}^{\ell}$ and $\mathcal{X}_{V}^{\ell+1}$
When the marked points are critical points, with $f:=f_{1}=\cdots=f_{m}$, Dujardin proved that there is a dense set of parameters in $\operatorname{supp}T_{1}\wedge\cdots\wedge T_{r}$ in $\mathcal{M}_{d}^{cm}$ at which the critical point ${\bf c}=(c_{1},\ldots,c_{r})$ for $(f,\ldots,f)$ is transversely prerepelling over $\mathcal{M}_{d}^{cm}$ [Du, Theorem 1.1]. And conversely, if the $r$-tuple of critical points is properly prerepelling for $(f,\ldots,f)$ at some parameter over a parameter space of dimension $\geq r$, then $T_{1}\wedge\cdots\wedge T_{r}$ must be nonzero in that parameter space, as first observed in [BE] (where it was proved in the case $r=2d-2$ and assuming transversely prerepelling) and revisited in [Ga1, Theorem 6.2] and [BBD, Proposition 3.7] in more general contexts and without assuming transversality. See [DM1, Proposition 4.8] for the version we use.
@@ 3.1. Measure on $\mathcal{X}_{V}^{\ell}$
If ${\bf c}$ satisfies the conclusion of Proposition 3.1 for $\Phi$, we say that it witnesses the nonvanishing of $\mu_{\ell}$.
@@ 3.2. Measures on $\mathcal{X}_{V}^{\ell+1}$
In what follows it is convenient to work with the Deligne pairing (of relative dimension dimension 00), of the invariant metrized line bundles restricted to graphs of critical marked points (see [YZ, §4.2] for metrics of the Deligne pairing and [YZ, §6.2.2] for the ‘restriction’ of the invariant metrized line bundle to a subvariety). By [YZ, Theorem 4.2.3], to each marked tuple of critical points ${\bf c}:\mathcal{X}^{\ell}_{V}\to(\mathbb{P}^{1})^{\dim\mathcal{X}^{\ell}_{V}}$ we can associate a metrized line bundle $\langle\overline{L}_{\Phi}|_{\Gamma_{\bf c}}\rangle\in\widehat{\mathrm{Pic}}(\mathcal{X}^{\ell}_{V})_{\mathbb{Q},\mathrm{nef}}$, where $\Gamma_{\bf c}$ denotes the graph of ${\bf c}$ in $\mathcal{X}^{\ell}_{V}\times(\mathbb{P}^{1})^{\dim\mathcal{X}^{\ell}_{V}}$. Moreover, we let
@@
and note that their curvature forms (at an archimedean place) are the bifurcation currents
@@
Recall that $T_{f_{i},c_{j}}^{\wedge 2}=0$ for all $i,j$ and similarly $(\Pi^{*}_{2}c_{1}(\overline{L}_{i}))^{\wedge r+1}=(\Pi_{2})^{*}(c_{1}(\overline{L}_{i})^{\wedge r+1})=0$ for $i=1,2$, by the definition of the bifurcation current as limits of pullbacks, because the map $\Phi_{(\ell+1)}$ acts on the $r$-dimensional $(\mathbb{P}^{1})^{r}$. Moreover, since $\dim\mathcal{X}^{\ell}_{V}=\ell(r+1)$ we also have that $\Pi_{1}^{*}(c_{1}(\overline{L}))^{\wedge\ell(r+1)+1}=0$. Thus , using also [YZ, Theorem 4.2.3] as before, we have
@@
for all $(\lambda,f_{1},\ldots,f_{\ell+1})\in U(\overline{\mathbb{Q}})$ for a Zariski open and dense subset $U\subset\mathcal{X}^{\ell+1}_{V}$.
It remains to show that for some $\epsilon=\epsilon(\mathcal{X})$, the set
@@
as measures on $X_{\lambda}$, for all $\lambda$ in the support of $\pi_{*}\mu_{\ell}$ in $V$. Indeed, suppose there exists $b\in V(\mathbb{C})$ where (4.8) doesn’t hold, and so that $\pi_{*}\mu_{\ell}(U)>0$ for every open neighborhood $U$ of $b$. Then we can find a continuous, non-negative, compactly supported function $\psi$ on $\mathcal{X}$ such that
@@
Recall from Proposition 3.2 that we have parameter $x_{1}=(\lambda_{1},s_{1},\ldots,s_{\ell})\in\operatorname{supp}\mu_{\ell}$ so that the tuple of marked critical points $(c_{1},\ldots,c_{r})$ is transversely prerepelling for $(f_{\ell},\ldots,f_{\ell})$ at $s_{\ell}$ along $X_{\lambda_{1}}$. This implies that $\mu_{\lambda_{1}}^{r+1}$ is nonzero on $X_{\lambda_{1}}$, with $s_{\ell}$ in its support. Moreover, as in [BE, Section 3] and the generalization appearing in [Ga2, Proposition 3.2], we know that if we zoom in on the measure $\mu^{r+1}_{\lambda_{1}}$ on $X_{\lambda_{1}}$ at the parameter $s_{\ell}$ at an exponential rate in $n$ (determined by the multipliers of the repelling cycles in the forward orbits of the marked critical points), and rescale the measure by $d^{nr}$, we get a nontrivial limit.
On the other hand, the critical point $c_{r+1}$ is attracted to a parabolic cycle for $f_{\ell}$ at the parameter $x_{1}$, by construction. By (4.8), we know that $\mu_{\lambda_{1}}^{1}=\alpha^{-1}\mu_{\lambda_{1}}^{r+1}\not=0$ on $X_{\lambda_{1}}$. But because of the parabolic point, we claim that if we zoom into the measure $\mu_{\lambda_{1}}^{1}$ on $X_{\lambda_{1}}$ at $s_{\ell}$ at the same exponential rate (which is supposed to give a nonzero limit), we actually get 0. This will produce our desired contradiction.
@@
Suppose $f_{x}$ is a holomorphic family of maps on $\mathbb{P}^{1}$ parametrized by $x\in\mathbb{D}^{r}\subset\mathbb{C}^{r}$, with a marked critical point $c:\mathbb{D}^{r}\to\mathbb{P}^{1}$. Assume that $c(0)$ is attracted to a parabolic fixed point at $z=0$ for $f_{0}$. Then for any constant $\rho>1$, there exists $\delta>0$ so that the sequence of functions